OpenStudy (anonymous):

Find the following product: (x+8)²

OpenStudy (anonymous):

i got xᶾ+64

OpenStudy (unklerhaukus):

\[(x+8)^2\]\[=(x+8)(x+8)\]\[=x(x+8)+8(x+8)\] \[=\]

OpenStudy (anonymous):

\[(a+b)^2 = a^2+2ab+b^2 \]